### v1

# Mixed Strategy Constraints in Continuous Games

Abstract.
Equilibrium problems representing interaction in physical environments typically require continuous strategies which satisfy opponent-dependent constraints, such as those modeling collision avoidance. However, as with finite games, mixed strategies are often desired, both from an equilibrium existence perspective as well as a competitive perspective. To that end, this work investigates a chance-constraint-based approach to coupled constraints in generalized Nash equilibrium problems which are solved over pure strategies and mixing weights simultaneously. We motivate these constraints in a discrete setting, placing them on tensor games ($n$-player bimatrix games) as a justifiable approach to handling the probabilistic nature of mixing. Then, we describe a numerical solution method for these chance constrained tensor games with simultaneous pure strategy optimization. Finally, using a modified pursuit-evasion game as a motivating examples, we demonstrate the actual behavior of this solution method in terms of its fidelity, parameter sensitivity, and efficiency.
conference: ; ;
1. Introduction
When solving Nash equilibrium problems for applications in robotics, control, and automation, there an assortment of practical considerations to be addressed. These games are typically played in physical space, and therefore are modeled with continuous strategy spaces. As with finite games, mixed strategies are often required for continuous games to ensure the existence of an equilibrium. However, physical-space games often also have constraints involving the strategies of multiple players (e.g. collision constraints: no two players can occupy the same space at the same time). Properly handling constraints in a mixed-strategy framework is the focus of this work.
A more simple application of these so-called coupling constraints occurs in discrete games with mixed strategies, before any explicit consideration of improving the pure strategies. Coupling constraints are troublesome in the presence of mixing, even in discrete settings: since mixed strategies are probabilistic, a constraint may or may not be violated depending on the strategies actually realized by opponents. We determine that chance constraints can be employed to intuitively account for mixing.
Of greater interest, however, are continuous-strategy games. In general, continuous games require mixing over an infinite support of pure strategies to ensure the existence of a Nash equilibrium. However, in certain classes of games such as separable games, finite support mixed strategy equilibria exist ((Stein et al., 2008)). We seek pure strategies that admit a competitive mix with low cost, the pure strategy space can itself be subject to optimization, as in (Peters et al., 2022). In practice, mixing is quite desirable in continuous games, often providing a competitive advantage even with minimal support.
In this paper, we place multiple-player constraints on mixed strategy games subject to continuous pure strategy optimization. We employ chance constraints in this continuous setting, describing in detail the necessary modifications required to function in a pure strategy optimization setting. Keeping in mind our goal of formulating games in physical space, we consider a modified, three-player constrained pursuit-evasion game. We conclude with promising areas of further investigation, and a discussion of the application of these constraints to model common problems in our spatial domain.
1.1. Our Contributions
This work presents the following contributions:
  • A formulation of chance constraints intended to be applied to mixed-strategy tensor games with tensor constraints, a bimatrix game generalization;
  • A numerical solver which uses this formulation to seek mixed-strategy Nash equilibria, optimizing along both the pure strategy space and the mixing weights;
  • Empirical behavior of the solver when applied to constrained pursuit-evasion games,
  • Analysis of the solver quality with respect to its parameters, particularly regarding the feasibility of the resulting solution.
The structure of this paper is as follows: In the remainder of this section we briefly discuss the related work on generalized Nash equilibrium problems and chance constraints to scope our work. In Section 2 we define tensor games with tensor constraints, a mathematical preliminary to working with chance constraints under pure strategy optimization, and in Section 3 we discuss those constraints in discrete settings. Section 4 covers our solver with which we compute the empirical results discussed in Section 5, followed by a brief consideration of future extensions and conclusion.
1.2. Related Work
1.2.1. Coupled constraints
For our purposes in constraining game-theoretic behavior in real space, we require a notion of constraints involving multiple players. These constraints are known as shared or coupled constraints, and games [^1] There is some nuance as to whether GNEPs are, in fact, classifiable as games. In physical-space problems where mutual interactions are necessary and their corresponding constraints are as natural as single-player constraints, this distinction is of little importance. which feature them are known as generalized Nash equilibrium problems, or GNEPs, as described in (Facchinei and Kanzow, 2010a). Other (older) terminology, like that seen in (Harker, 1991), refers to coupled constraints as feasible set interactions, found in social equilibrium problems and pseudo-Nash games.
Solutions to solving GNEPs vary, and generally do not use an explicit concept of mixing. (Krawczyk, 2007) summarizes numerical solutions to solving these games, a project further advanced in, e.g., (von Heusinger, 2009). Past avenues include penalty methods (Facchinei and Kanzow, 2010b), augmented Lagrangian methods (Kanzow and Steck, 2016), and methods informed by error bound theory (Ba and Pang, 2022). This work focuses on the representation and modeling of constraints in mixed-strategy games, and less on computation. In this work we characterize GNEPs as a mixed complementarity problem via the KKT conditions, and solve them via a readily available solver (Dirkse and Ferris, 1995).
Our primary divergence from these works lies in the presence of pure strategy optimization, where we consider both pure strategies and their mixing weights in finding a Nash equilibrium. To our knowledge, this specific form of generalized Nash equilibrium has never been fully explored in the literature, which prevents further development of mixed-strategy game-theoretic agents in constrained physical-space games.
1.2.2. Chance constraints
To intuitively account for the stochasticity of mixed strategies, we use chance constraints, originally formulated in (Charnes and Cooper, 1959). Chance-constrained problems without feasible set interactions (that is, in ordinary Nash equilibrium problems) are well covered, and significant work has been done to characterize the circumstances in which such problems have equilibria (as in (Henrion and Strugarek, 2008), (Geletu et al., 2013)). Chance constraints have been used in the presence of mixing for decades, see e.g. (Mukherjee, 1980).
Some work has been done describing the behavior of chance constraints in GNEPs, but never with the intent of pure strategy optimization. (Zhang et al., 2022) manage mixing chance constraints in discrete mixed-strategy scenarios from the variational inequality point of view, while (Singh and Lisser, 2018) apply the same treatment to a continuous single strategy to handle stochastic payoffs, but the former assumes given pure strategies, while the latter does not mix. We assume pure strategies as additional, continuous decision variables rather than as given, and mix over them under the presence of chance constraints.
2. Preliminaries: Tensor Games with Tensor Constraints
We begin with then necessary mathematical preliminaries to discuss (generalized) Nash equilibrium problems under pure strategy optimization when all players interact in constraints.
In an $n$-player game, the resulting expected cost is a tensor, with each player mixing along each axis of the tensor. As such we consider tensor games, (sometimes referred to as multilinear games), $N$-player extensions of bimatrix games. Such games are established in the literature ((Papadimitriou and Roughgarden, 2005), (Huang and Qi, 2017)); they are known to have mixed-strategy equilibria and multiple solution methods exist. We specifically formulate tensor games with tensor constraints, allowing all players to affect all constraints through their mixing probabilities or, in continuous games, through the cost and constraint tensors.
2.1. Definitions
Tensor games occur when each player $i$ mixes over $m_{i}$ pure strategies with probabilities $x_{i}$. We denote player $i$’s $k$-th pure strategy $s_{i,k}$ and that player’s pure strategy space as $\mathcal{S}_{i}$, such that $s_{i}\in\mathcal{S}_{i}^{m_{i}}$ (and similarly denoting $s\in\mathcal{S}$).
In tensor games, each player has a cost tensor; each element of the cost tensor corresponds to a joint strategy. Cost tensors are denoted $A_{i}$, and all $A_{i}\in\mathbb{R}^{m_{1},..,m_{i}}$. The resulting cost, when each player has probability $x_{i,k}$ assigned to their $k$th strategy, is a homogeneous $N$-polynomial. For ease, we denote the operation that forms this polynomial $A_{i}\;[\cdot]\;x$, under the following definition:
Definition 2.1 ().
For a tensor $T$ and vectors $x_{1}..x_{N}$, $T\;[\cdot]\;x$ denotes the dot product of every vector $x_{i}$ along the $i$th axis of $T$. That is,
(1) $T\;[\cdot]\;x:=(((T\cdot x_{i})\cdot..)\cdot x_{N}$
We also constrain tensor games with a total of $C$ constraints (indexed by $j)$. Each player may or may not respect a particular constraint; we denote the index set of constraints respected by player $i$ as $\mathcal{O}_{i}$. Each constraint tensor is denoted as $Q_{j}$ and has the same shape as the cost tensor. The elements of the constraint tensor represent some value regarding the feasibility of the joint strategy with the corresponding index, and we constrain the expected value: $Q_{j}\;[\cdot]\;x\geq\epsilon_{j}$ for some arbitrary $\epsilon_{j}$.
With this notation in place, we define tensor games with tensor constraints:
Definition 2.2 ().
A tensor game with tensor constraints $G(A,Q,\epsilon)$ takes the following form, given the cost tensors $A_{i}$, constraint tensors $Q_{i,j}$, and thresholds $\epsilon_{i,j}$:
(2) $\forall i\in\{1..N\}:\begin{bmatrix}\;\;\min\limits_{x_{i}\in\Delta^{m_{i-1}}}\;\;A_{i}\;[\cdot]\;x\;\;\;\;\textrm{s.t.}\;\;\;\;\\[8.0pt] \;\;\;\;\;\;Q_{i,j}\;[\cdot]\;x-\epsilon_{i,j}\geq 0\;\;\forall j\in\mathcal{O}_{i}\end{bmatrix}$
Let $x_{\neg i}$ denote the concatenation of strategy vectors $x_{1},...,x_{i-1},x_{i+1},...,x_{N}$, and let $(x_{i},x_{\neg i})$ be shorthand for $x$.
Definition 2.3 ().
An equilibrium point for a game $G(A,Q,\epsilon)$ is a strategy vector $x^{*}$ such that for all $i\in\{1,...,N\}$,
(3) $A_{i}\;[\cdot]\;(x_{i}^{*},x_{\neg i}^{*})\leq A_{i}\;[\cdot]\;(x_{i},x_{\neg i}^{*})\ \forall x_{i}\in\Delta^{m_{i-1}}:Q_{i,j}\;[\cdot]\;(x_{i},x_{\neg i}^{*})-\epsilon_{i,j}\geq 0\ \forall j\in\mathcal{O}_{i}$
In discrete games, the way in which the contents of the cost and constraint tensors are derived is external to the game. However, for our eventual purpose of applying constraints on continuous mixed-strategy games to model physical-space interactions, we assume that the elements of the tensors are related the pure strategies by way of real-valued cost and constraint functions $f_{i}$ and $g_{i,j}:\mathcal{S}\rightarrow\mathbb{R}$. Let $K$ be an index into a cost or constraint tensor (and therefore $K_{i}$ is the index of the strategy realized by player $i$ in a particular joint strategy). Then, the cost and constraint tensor elements are defined accordingly:
(4) $A_{i,K}=f_{i}(s_{1,K_{1}}\;..\;s_{N,K_{N}})\;\;;\;\;Q_{i,j,K}=\rho(g_{i,j}(s_{1,K_{1}}\;..\;s_{N,K_{N}}))$
We use the convention that $g_{i,j}(s_{K})\geq 0$ indicates a satisfied constraint (that is, a feasible joint strategy).
When using pure strategy optimization, our goal (from the perspective of player $i$) is to both mix over pure strategies correctly, and to select $s_{i}$ that generate the most favorable $A$ and $Q$ under a correct mix. For this purpose we assume $f$ and $g$ are twice differentiable and defined over a compact $\mathcal{S}$.
The function $\rho_{j}$ is the constraint indicating function. In theory, $\rho(f(s_{K}))=\text{1}(f(s_{K})\geq 0)$, but we will consider examples where this is not the case.
2.2. Existence of equilibria
Tensor games with tensor constraints do not always admit and equilibrium point. This can cause problems for the optimization of pure strategies. As such, we briefly pause to establish sufficient conditions to ensure the presence of an equilibrium in a tensor-constrained tensor game.
Theorem 2.4 ().
Tensor games with tensor constraints have at least one equilibrium if, for all sets of opponent strategies $x_{\neg i}$, there is a feasible ego strategy $x_{i}$, and an open set of feasible strategies around it.
Proof.
Using the notation from Eq. 2 and Eq. 4, let $C_{i}(x_{\neg i}):=\{x_{i}:\forall j\in\mathcal{O}_{i}:Q_{i,j}\;[\cdot]\;x-\epsilon_{i,j}\geq 0\}$ denote the constraint correspondance from joint opponent strategies to the set of feasible ego strategies. Let $\theta(x):=A^{(i)}\;[\cdot]\;x$ denote the overall cost function for each player.
We then follow the conditions put forth by (Dutang, 2013):
  1. (1)
    $\Delta^{m_{i}-1}$, a simplex, is a non-empty, compact, convex polytope.
  2. (2)
    $A_{i}\;[\cdot]\;x$ is continuous on graph $\textrm{Gr}(C_{i}(x_{\neg i}))$ because it is a multilinear map, which is continuous on the entire joint strategy space.
  3. (3)
    $x_{i}\rightarrow\theta_{i}(x_{i};x_{\neg i})$ is quasiconcave on $C_{i}(x_{\neg i})$ for all $x_{i}\in\Delta^{m_{i}-1}$ because $\theta_{i}$ is a multilinear map and therefore linear given $x_{\neg i}$.
  4. (4)
    $C_{i}(x_{\neg i})$ is nonempty, closed and convex for all $x_{\neg i}\in\Delta^{m_{\neg i}-1}$. This follows from the assumption that each player has an open set of feasible strategies given any opponent strategies.
  5. (5)
    First, $Q_{j}\;[\cdot]\;x$ is continuous and concave in $x_{i}$ given all $x_{\neg i}$, since it is linear in that case. Second, we have assumed that $C_{i}(x_{\neg i})$ contains some nonempty open set of feasible solutions. These two facts satisfy Prop. 4.2 of (Dutang, 2013), which guarantees lower and upper semicontinuity of $C_{i}$ in $x_{\neg i}$
With these conditions satisfied, tensor games with tensor constraints satisfying the assumption must have at least one equilibrium. ∎
Note that the open-set assumption is sufficient, but not necessary; there are tensor games with equilibria that do not satisfy it.
2.3. As mixed complementarity problems
Since the optimization problems in (2) are all linear in the decision variable $x_{i}$, tensor games with tensor constraints can exactly be reformulated as mixed complementarity problems by way of the KKT conditions; we use this approach here to solve them. Note that in a bimatrix game, the corresponding complementarity problem is linear, but in this is not necessarily true in $n$-player tensor games, which have multilinear terms. The full mixed complementarity problem is as follows:
(5) $\nabla_{x_{i}}L_{i}(x,\lambda,\gamma)$ $\perp 0\leq x_{i}\leq\infty,$
$\sum_{k=1}^{m_{i}}x_{i,k}-1$ $\perp-\infty\leq\lambda_{i}\leq\infty,$
$Q_{i,j}\;[\cdot]\;x-\epsilon_{i,j}$ $\perp 0\leq\gamma_{i,j}\leq\infty,\ \forall j\in\mathcal{O}_{i}$
with the Lagrangian function
(6) $L_{i}(x,\lambda,\gamma)=A_{i}\;[\cdot]\;x-\lambda_{i}\sum_{k=1}^{m_{i}}x_{i,k}-\sum_{j\in\mathcal{O}_{i}}\gamma_{i,j}(Q_{j}\;[\cdot]\;x-\epsilon_{j})$
for all $i\in 1..n$ and $j\in\mathcal{O}_{i}$. The complementarity notation $F_{i}(Z)\perp l_{i}\leq Z_{i}\leq u_{i}$ means either $F_{i}(Z)\geq 0$ and $Z_{i}=l_{i}$, or $F_{i}(Z)=0$ and $l_{i}\leq Z_{i}\leq u_{i}$, or $F_{i}(Z)\leq 0$ and $Z_{i}=u_{i}$. The second complementarity condition in Eq. 5 is simply an equality ensuring that the mixing weights sum to 1; the corresponding dual $\lambda_{i}$ is free. Here, $\lambda_{i}$ are duals for the (implied) constraint that probabilities sum to 1, and $\gamma_{i,j}$ are duals for the tensor constraints.
3. Discrete Multiple-Player Constraints
With our necessary preliminaries complete, but before discussing pure strategy optimization, we pause to conceptualize tensor constraints as chance constraints to account for the probabilistic nature of mixing.
3.1. Chance Constrained Tensor Games
3.1.1. Definition
When every element of $Q_{i,j}$ is the Boolean feasibility of the corresponding joint strategy (1 if the joint strategy is feasible, 0 if it is not), $Q_{i,j}\;[\cdot]\;x$ is the probability with which player $i$’s constraint $j$ is satisfied (at any value). This is an intuitive way to constrain a mixed game, where the feasibility of the realized joint strategy is indeed probabilistic. The value $\epsilon_{i,j}$ in Eq. 2 has the interpretation of a “confidence”, the minimum acceptable probability of an infeasible solution being realized.
To ensure a Boolean constraint tensor when a constraint function is available, $\rho(g(s_{K}))=\text{1}(g(s_{K})\geq 0)$ may be used. In practice we will find that an approximation to this is valuable.
3.1.2. Properties
The primary appeal of chance constraints comes from the ability to vary $\epsilon$ to soften constraints and remove the undesirable dependence of the solution on arbitrarily poor-quality, but feasible, pure strategies. As such, chance constraints are applicable in “dense” problems with many players and significant mixing, in which many joint strategies have low feasibility due to the high number of player interactions. The existence of a solution is maintained by allowing players to mix with all strategies, not only those with no possibility of infeasibility; as the number of pure strategies grows the possibility of infeasibility disappears. In this way we might expect the general amount of mixing to decrease as $\epsilon$ increases.
On the other hand, $\epsilon$ also has an effect on the solution quality; low settings of it can permit low-feasibility equilibria that are filtered out at higher confidences. In games that have mixed strategy equilibria where the constraint is fully upheld, we might expect the amount of mixing to increase with $\epsilon$.
At $\epsilon_{i,j}=1$, all combinations of pure strategies must be feasible w.r.t. the constraint function $g_{i,j}$, which can lead to unusual behavior. As an example, consider a game where each of two players has $m$ strategies and is incentivized to mix over all of them, with the caveat that the two players cannot choose the same strategy. At $\epsilon_{i,j}=1$, as $m\rightarrow\infty$, it is impossible for all players to achieve an even mix for any finite but arbitrarily large number of lanes; there is always a (vanishingly small) likelihood of collision.
3.2. An Alternative: Constraints on Expectation
There is a point of view from which chance constraints appear needlessly verbose, particularly as we develop the machinery needed to apply them to pure strategy optimization. Why is the constraint indicating function $\rho$ necessary at all - why not define the $K$-th element of $Q_{i,j,K}$ directly as $g_{i,j}(s_{K})$, in the same manner we define the elements of the cost tensor $A$?
This is interpretable as constraining the expectation of the feasibility, and in fact, expectation constraints are somewhat justifiable in discrete games. They have a well-defined, intuitive meaning when the pure strategies are limited in number and static. Also, when using penalty methods (under a scalar penalty, perhaps a dynamic one), expectation-constrained games become ordinary tensor games with no irregular behavior, aside from that inherited from $f$ and $g$.
However, expectation constraints have an unbounded dependence on the constraint function and also, (depending on that constraint function) on the strategies. As we shall see, the purpose of $\rho$ is to bound this dependence. Otherwise, it is possible for changes in a pure strategy to cause arbitrarily large changes in feasibility. That is, if some pure strategy $s_{i,k}$ causes an extremely high feasibility, all other strategies owned by player $i$ have no comparative effect on the feasibility. An arbitrarily small mixing weight (inversely small with respect to the feasibility value) can be assigned to $s_{i,k}$ and it is as though the constraint is not present.
We seek to avoid this behavior when using pure strategy optimization, and we will return to this example to determine if this poor case in fact occurs. First, however, we develop a solver capable of optimizing pure strategies in tensor games with tensor constraints in the following section.
4. Implementing Constrained Pure Strategy Optimization
Thus far, we have discussed tensor games with tensor constraints, where the tensor constraints capture feasible set interactions in the presence of strategic mixing. We now proceed to describe our method of performing pure strategy optimization, i.e. choosing both the pure strategies that result in cost and constraint tensors, as well as the mixing weights over these tensors.
As an aside, the method of pure strategy optimization we suggest here is relatively simple: we append the pure strategies to the decision variables of the tensor game, formulate a mixed complementarity problem using the KKT conditions, and solve. However, note that there are other ways of performing pure strategy optimization; for instance, (Peters et al., 2022) formulate it as a two-level optimization by calculating gradients through an ordinary tensor game. We find the simple approach to be faster for the problems we explore in this work, although other approaches an interesting topic for future research.
4.1. Strategy-Augmented Tensor Games with Tensor Constraints
To model equilibrium seeking in both the strategy space and the weight space, the simplest approach is to include the pure strategies as decision variables of the GNEP. This results in a variation on a tensor game with tensor constraints, with the following form:
[…]
Under differentiable $f$, $g$, and $\rho$, the KKT conditions of this formulation can be taken, forming a new mixed complementarity problem where the cost and constraint tensors are functions of the strategies:
(8) $\nabla_{x_{i}}L_{i}(x,s,\lambda,\gamma)$ $\perp 0\leq x_{i}\leq\infty$
$\nabla_{s_{i}}L_{i}(x,s,\lambda,\gamma)$ $\perp-\infty\leq s_{i}\leq\infty$
$\sum_{k=1}^{m_{i}}x_{i,k}-1$ $\perp-\infty\leq\lambda_{i}\leq\infty$
$Q_{i,j}(s)\;[\cdot]\;x-\epsilon_{i,j}$ $\perp 0\leq\gamma_{i,j}\leq\infty,\ \forall j\in\mathcal{O}_{i},$
with the Lagrangian function:
(9) $L_{i}(x,s,\lambda,\gamma)=A_{i}(s)\;[\cdot]\;x-\lambda_{i}\sum_{k=1}^{m_{i}}x_{i,k}-\sum_{j\in\mathcal{O}_{i}}\gamma_{i,j}(Q_{i,j}(s)\;[\cdot]\;x).$
Unlike in the finite game setting, the optimization problems (7) are generally non-convex with respect to the decision variables. Therefore a solution to the joint KKT conditions (8) is not necessarily a proper equilibrium point. Nevertheless, for our purposes of exploring ways to model constraints in mixed-strategy continuous games, the KKT approach is satisfactory. The mixed complementarity problem implied by (8) can be solved with an ordinary MCP solver in the typical way. The difficulty of solving this MCP depends on the nonlinearity of $f$, $g$, and $\rho$; under linear functions the problem is multilinear and therefore comparable to the MCPs defined by (5).
4.2. Iterative Tightening
The approach above only holds under twice differentiable cost, constraint, and constraint indicating functions. While it is reasonable to expect this fact about the cost and constraint functions in spatial games, it does not hold for $\rho$ under default chance constraints, because the indicator 1 is non-differentiable. To circumvent this, we “soften” the indicator into a sigmoid function, so that
(10) $\rho(g(s))=\sigma(\omega g(s))=\frac{1}{1+e^{-\omega g(s)}},$
where $\omega$ sets the extent of softening. We seek to use the highest value of $\omega$ that is numerically feasible, since $\lim_{\omega\rightarrow\infty}\sigma(\omega x)=\text{1}(x\geq 0)$. If $\omega$ is too small, the function $\rho$ does not act as a good approximation to the indicator function, and the resulting tensor constraints will not be representative of proper chance constraints. If $\omega$ is too high, the constraint gradient will be lost numerically and the MCP will generally suffer from poor conditioning. Luckily, near the constraint boundary, this conditioning issue diminishes.
Using that insight, we apply a simple iterative tightening procedure. We first solve the problem with a low value of $\omega$; then, we double the strictness and solve the problem again, using the previous solution as an initialization. We repeat this process until we reach a desired strictness or the solver fails.
Algorithm 1 gives the overall approach, where solve_gnep($Q,A,\epsilon$) formulates the KKT conditions in (8) and solves the resulting MCP.
Input: Twice differentiable cost and constraint functions $f_{i}$, $g_{j}$; initial guesses $x$, $s$
Output: Optimal strategies and mixing weights $s$ and $x$ at equilibria among the players indexed by $i$
for indices $K$ in cost tensors $A_{i}$ do
      $A_{i,K}\leftarrow f_{i}(s_{K})$ ;
      end for
     repeat
           for indices $K$ in constraint tensors $Q_{j}$ do
                $Q_{j,K}\leftarrow 1/(1+\exp(-\omega\cdot g_{j}(s_{K})))$ ;
                end for
               $x$, $s$leftarrow solve_gnep($A$, $Q,\epsilon$; init=($x$, $s$));
                $\omega\leftarrow 2\omega$;
                until $\omega\geq 2\omega_{\textrm{des}}$;
ALGORITHM 1 Constrained Strategy/Weight Equilibrium Computation
4.3. Technical Details
4.3.1. Efficiency
Tensor games in general do not scale well in the number of players, as the cost tensors are exponential in size with $N$. This is a known problem studied in the literature, and there are a variety of methods to ameliorate it, from sampling methods (Gemp et al., 2021) to parallelization (Widger and Grosu, 2009) to neural network approaches (Marris et al., 2022). Fortunately, the addition of constraints does not further affect this complexity, as each constraint tensor is exactly the size of the cost tensors (and are often more sparse); scaling is linear with the number of constraints.
For clarity, in this paper, we assume only proper handling of sparsity in tensors, and otherwise use no efficiency modifications.
4.3.2. Tooling
We implement Algorithm 1 in Julia (Bezanson et al., 2017), a general-purpose mathematical programming language. We use the PATH solver ( (Dirkse and Ferris, 1995)) to find solutions to the underlying MCPs.
5. Evaluating Constraints under Pure Strategy Optimization
With an adequate solver for seeking equilibria in both continuous pure strategies and their weights under constraints, we now turn to characterizing its applicability towards spatial games with intuitive collision constraints. We primarily consider “pursuit-evasion” games. Pursuit-evader games are a useful benchmark in our considerations as in pure strategy dimension of one or higher, the evader is incentivized to mix in apparently “deceptive” behavior, while pure strategies themselves must be placed strategically as well (Peters et al., 2022).
5.1. Solution Fidelity and Parameter Settings
We begin by considering the adequacy of the solver for seeking constrained equilibria in pure strategies and weights with respect to its parameters.
5.1.1. The “Hog-Poacher-Ranger” game
Consider a constrained pursuit-evasion game for use in our discussions of solver fidelity. In this “hunt-in-a-hunt” game, three players (the “hog”, the “poacher”, and the “ranger”, indexed in that order) interact in a two-dimensional space:
  • The “hog” maximizes its squared distance from the “poacher”.
  • The “poacher” minimizes its squared distance to the “hog”, under the constraint that it must not be within a certain distance $r$ of the “ranger”.
  • The “ranger” minimizes its squared distance to the “poacher”.
This is a tensor game with a single tensor constraint (respected only by the “poacher” player). Each player mixes over $m_{i}$ pure strategies $s_{i}$, which are also to be selected optimally. Each strategy $s_{i,k}\in[-2,2]^{2}$. The cost and constraint functions are as follows, where $\hat{s_{i}}$ is the strategy realized by player $i$:
(11) $f_{1}(\hat{s})$ $=-||\hat{s}_{1}-\hat{s}_{2}||_{2}^{2}$
$f_{2}(\hat{s})$ $=||\hat{s}_{1}-\hat{s}_{2}||_{2}^{2}$
$f_{3}(\hat{s})$ $=||\hat{s}_{2}-\hat{s}_{3}||_{2}^{2}$
$g_{1}(\hat{s})$ $=||\hat{s}_{2}-\hat{s}_{3}||_{2}^{2}-r^{2}$
Figure 1. Potential (non-exhaustive) Nash equilibria of the Hog-Poacher-Ranger game with two pure strategies per player. Strategy weight is mapped to opacity. In (b), the “hog” player is rendered underneath the “ranger” player.
The hog-poacher-ranger game has multiple optimal settings of the pure strategies $s$ (that is, multiple Nash equilibria in $x$ and $s$). In general, mixed strategies are used by all players except the “ranger.” In the remainder of this paper we allot all players of this game two pure strategies over which to mix (regardless of whether they actually must do so at equilibrium) and use $r=1$.
5.1.2. Solving Hog-Poacher-Ranger
We deployed an implementation of the solver described in Alg. 1 on one thousand randomized initial solutions of the “hog-poacher-ranger” game for several parameter variations of chance constraints. We noted cases where the game was unable to be solved, as well as the average actual chance of violating the constraint, which generally differs from the constraint as computed during solving. We also note the cost of every player (note that the game is zero-sum between the “hog” and “poacher”) and the time taken to solve. Table 1 summarizes these results. We also include the behavior of expectation constraints as described in Section 3.2
Chance; ϵ=italic-$\epsilon=$50% Chance; ϵ=italic-$\epsilon=$80% Chance; ϵ=italic-$\epsilon=$99% Expectation
% Solved 90.70% 83.0% 73.1% 98.50%
Hog (neg. Poacher) Cost -3.261 -3.003 -4.319 -3.251
Ranger Cost 1.029 1.504 2.585 1.000
Feasibility Chance 44.04% 100.00% 100.00% 21.04%
Avg Solver Time (ms) 12.14 18.68 15.61 5.77
Table 1. Bulk results of playing the “hog-poacher-ranger” game. “Feasibility Chance” is the likelihood that the constraint is upheld after solving, applying the exact constraint to the solution.
5.1.3. Computation Results: N-Player Hog-Poacher-Ranger Game
To benchmark our solver, we also used an $N$-player variant of this game, in which the first player acts as the hog, the $N-2$ subsequent players act as poachers aiming to minimize expected squared distance to their predecessor, and constrained to maintain a minimum distance from their successor. The final player also aims to minimize the expected squared distance to its predecessor. Because this game is well-defined for $N$ players, we can perform a numerical analysis to evaluate the solution time when solving using Algorithm 1 for various values of $N$ as well as various values of $m$, where $m$ is the number of pure strategies held by each of the players. The results of this experiment are listed in the following table.
$m=2$ 3 4 5
N=2 0.0034±0.0005 0.0088±0.0003 0.0092±0.0004 0.0192±0.0006
3 0.0159±0.0007 0.0531±0.001 0.0521±0.0011 0.2213±0.003
4 0.0484±0.0092 0.0692±0.0008 0.2514±0.0025 0.2376±0.001
5 0.0596±0.0011 0.4768±0.0067 0.3692±0.0043 0.594±0.0018
6 0.0946±0.0022 0.4188±0.0008 0.6166±0.0027 1.3409±0.0024
7 0.0847±0.0015 0.9942±0.0017 1.198±0.006 2.8395±0.0149
Table 2. Solve times for the $N$-player variant of the Hog-Poacher-Ranger game, with $m$ actions for each player. Trials were evaluated over $10$ solves per configuration. Average times are shown in seconds, along with $95\%$ confidence estimates coming from the standard error.
5.1.4. Parameters: Confidence
Our application of chance constraints uses two types of parameter: $\epsilon_{i,j}$ and $\omega$. We explore the impact they have on the resulting game, considering first $\epsilon_{i,j}$. Figure 2 shows the actual average number of times the constraint was upheld in the “hog-poacher-ranger” game with respect to different settings of $\epsilon_{i,j}$ (that is, the percentage of games in which the poacher “escapes”). We averaged across one hundred games per setting and $\omega_{\textrm{des}}=64$ was used; games in which $\omega<10$ after iterative tightening were considered solver failures. The dotted line corresponds to the expected percentage of “escapes” (the exact constraint). We might expect this to become equal to $\epsilon_{i,j}$, but in fact, this only occurs at $\epsilon\in\{0,0.5,1.0\}$. We see that settings above $\epsilon_{i,j}=0.5$ the solution tends to be overly conservative, while below $\epsilon_{i,j}=0.5$, the solution tends to be loose.
Figure 2. Feasibility at equilibrium for solves of the ordinary “hog-poacher-ranger” game using Alg. 1.
We can rationalize this behavior as being caused by the limited number of mixed strategies available. In this problem, at any equilibrium, the poacher mixes their two strategies with weights [0.5, 0.5] and the ranger does not mix. Therefore, a feasibility which is not in $\{0.0,0.5,1.0\}$ is impossible with a perfectly strict constraint (as the poacher can mix zero, one, or two strategies). However, during solving, only the soft constraint is available, and $\epsilon$ determines where strategies lie on it. At $\epsilon<0.5$, pure strategies may lie slightly inside (violating) the true constraint; at $\epsilon>0.5$ they must lie slightly outside it. When the perfectly strict version of the constraint is then used these two cases appear to have opposite feasibility, even when they may both lie very close to the constraint boundary.
5.1.5. Parameters: Strictness
Based on the behavior of $\epsilon$ we expect $\omega$ to bring the pure strategies closer to the true constraint as it is increased. We investigate this further by investigating each step of the iterative tightening procedure. For each of one hundred initial positions per setting of $\epsilon$, we attempt Alg. 1 on each with $\omega_{\textrm{des}}=64$. We track the minimum realizable distance between the “poacher” an the “ranger” as a proxy to the fidelity with which the constraint is upheld (or overshot) for each $\omega$.
Figure 3. Effect $\omega$ on constraint value (minimal distance between “ranger” and “poacher”). Actual constraint at distance of 1.
Figure 3 shows the results. As we should expect, as $\omega$ increases, the constraint indeed becomes closer to the constraint, and the extent to which this occurs depends on the severity of $\epsilon$. When $\epsilon<0.5$, the pure strategies converge to the constraint boundary from the inside of the constraint, which explains the unexpected relationship between confidence and true feasibility showcased in Fig. 2.
5.2. Outlier Sensitivity of Expectation Constraints
As expected, the behavior explained in Section 3.2 causes poor performance: selecting one extreme-feasibility strategy with low probability allows low-cost, infeasible strategies to be used with high probability. When pure strategies are continuous and subject to optimization, it is always possible to find such a strategy, exacerbating the problem in this setting.
For instance, in the “hog-poacher-ranger” example, the poacher may select pure strategies from the space $[-2,2]^{2}$, and in fact often selects a pure strategy on the border that allows the “primary” strategy to violate the constraint. Figure 4 showcases this common behavior when using expectation-based constraints.
Figure 4. One solution of the “hog-poacher-ranger” game under expectation constraints. Strategy weight is mapped to opacity.
This is not, strictly speaking, a solver failure, but rather a modeling mistake: the expected poacher-ranger distance correctly lies against the constraint (expected distance of 1). In fact, since expectation constraints are generally well-conditioned and no softening or additional parameters are required, solving this formulation is quick and rarely fails. However, expectation constraints defy our notion of “constraints”; they provide us no guarantees on the behavior under any particular realized strategies. We would like to be able to weigh between the feasibility of every particular pure strategy, and the feasibility of the strategy set as a whole, with some parameter of the constraint.
5.3. Promotion of mixed strategies
In Section 3.1.2, we surmised that chance constraints with a higher confidence could either prevent mixing (on the basis that most joint strategies can be infeasible and mixes are less likely to satisfy the constraint under high confidence), or promote it (on the basis that games which really do have a completely feasible mixed strategy solution are more likely to find it.) To build intuition on this matter for continuous games, we plot the actual “poacher” equilibrium weights achieved for each of one hundred trials across settings of $\epsilon$ in Figure 5.
Figure 5. Mixing weights of the “poacher” player for each constraint confidence trial in the “hog-poacher-ranger” game. (The horizontal axis is discrete; points are arranged laterally for each $\epsilon$ for visibility.)
We find that the “hog-poacher-ranger” game typically falls in the latter category: its equilibria are generally perfectly mixed for the “poacher” player, varying from this only at sufficiently low $\epsilon_{i,j}$. This is reasonable, based on the equilibria of this game shown in Figure 1; the true equilibria seem to have equal mixing for the poacher and we should expect to approach that as $\rho$ approaches the indicator 1.
In general, correlating the findings of Figs. 5 and 2, we find that the “hog-poacher-ranger” game is not sensitive to settings of $\epsilon$ given a sufficient $\omega$; all equilibrium strategies occur at perfect poacher mixing. However, this is not always the case, particularly as the number of mixed strategies increases. For instance, in a version of the “hog-poacher-ranger” game where the poacher is allowed to mix over three strategies, at $\epsilon<1$ equilibria exist where the poacher player overlaps the ranger player with probability $1-\epsilon$ and mixes equally over the remaining two strategies. Table 3 shows this relationship between $\epsilon$ and the realized feasibility as computed on a simplified “hog-poacher-ranger“ variant where the “hog” is stationary at (0, 0) and the “poacher” has three pure strategies to mix over.
Confidence ($\epsilon$) 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Realized Feasibility 0.1992 0.4043 0.5901 0.6528 0.7013 0.7536 0.8062 0.8797
Table 3. Relationship between the confidence value $\epsilon$ and the realized (exact) feasibility at equilibrium for a modified “hog-poacher-ranger”. Averages of twenty solves per confidence; $\omega_{\textrm{des}}=64$.
As we would expect, here, we do not see an immediate shift in realized feasibility at $\epsilon=0.5$ as we do in Fig. 2. Now, with one pure strategy not on the constraint boundary, the confidence $\epsilon$ becomes relevant to mixing. This is both a feature and a drawback of chance constraints: When many mixed strategies are present, this effect of $\epsilon$ helps prevent low-probability strategies from interfering and preventing the game from being solved, but doing so can (as in this case) obscure the true, exactly constrained equilibria, if there are any.
6. Closing Remarks
We end with some brief commentary about important modeling aspects when applying chance constraints under pure strategy optimization not covered here.
6.1. Extensions
6.1.1. Fairness
In general, tensor games with tensor constraints can admit multiple solutions. This can occur the same constraint is repeated under different ownership - for instance, $\mathcal{O}_{1}=\mathcal{O}_{2}$, $Q_{1,j}=Q_{2,j}$. This is generally undesirable in problems in physical space (particularly with regards to collision constraints in which, loosely speaking, we expect all players to “push” equally against the constraints). Ideally we seek a fair equilibria (Rabin, 1993).
When constraints appear in the same ownership sets, the formulation used here in fact yields a refinement of generalized Nash equilibria known as the variational equilibrium. The variational equilibrium is a sort of “fair” equilibrium, in that players share dual variables $\gamma_{i,j}$ for any particular constraint (as in (Schiro et al., 2013) and (Nabetani et al., 2011)). Unfortunately, this refinement is not sufficient to ensure a unique solution for tensor games with tensor constraints.
Here, we ignore this characteristic, assuming is generally the case that $x_{i}>0$. Nevertheless, it is important to remain aware of this possible redundancy to avoid unexpected complications during further optimization (i.e., over pure strategies). Furthermore, optimal pure strategy selections are generally not unique either, implying that there may be further-reaching fairness effects beyond the mixing weights alone.
6.1.2. Information Structure
Chance constraints can be applied to tensor games in bilevel structures. In this situation, upper-level (leader) players can intentionally spread their pure strategies and mix more evenly to block lower-level players using these constraints. Furthermore, the confidence $\epsilon$ is compelling as a knob for controlling the extent to which the upper level player may control the lower level through the constraint. As information structure becomes complicated beyond bilevel problems, these dynamics could prove quite nuanced indeed, calling for extensive tooling to handle feasible set interactions numerically.
6.2. Conclusion
We have presented a method for constraining equilibrium problems taking place in both pure strategies and mixing weights using a chance constraint based method. From the point of view of Nash equilibrium users - specifically roboticists, control theorists, and motion planners - the ability to handle feasible set interactions without sacrificing either strategic mixing or pure strategy optimality is critical to modeling problems in physical space. We look forward to applying this work to more complex spatial games and anticipate its utility in overcoming a recurring frustration in modeling agent interactions in game theoretic settings.